The following table displays the mathematics test scores for a random sample of, along with their final SY16C grades.

a.Fit the regression line y=a+bx to the data

and interpret the results.

b.Use the regression equation to determine the SY16C grade for a college student who scored 60 on their achievement test. What would their SY16C grade be?


Mathematics (x)SY16C grade (y)139652437832152464825579264789728738759893456




1
Expert's answer
2021-12-06T12:23:39-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 39 & 65 & 2535 & 1521 & 4225 \\
 \hdashline
 & 43 & 78 & 3354 & 1849 & 6084 \\
 \hdashline
 & 21 & 52 & 1092 & 441 & 2704 \\
 \hdashline
 & 64 & 82 & 5248 & 4096 & 6724 \\
 \hdashline
 & 57 & 92 & 5244 & 3249 & 8464 \\
 \hdashline
 & 47 & 89 & 4183 & 2209 & 7921 \\
 \hdashline
 & 28 & 73 & 2044 & 784 & 5329 \\
 \hdashline
 & 75 & 98 & 7350 & 5625 & 9604 \\
 \hdashline
 & 34 & 56 & 1904 & 1156 & 3136 \\
 \hdashline
 Sum= & 408 & 685 & 32954 & 20930 & 54191 \\
 \hdashline
\end{array}"



"\bar{X}=\dfrac{1}{n}\sum_iX_i=\dfrac{408}{9}\approx45.333333"

"\bar{Y}=\dfrac{1}{n}\sum_iY_i=\dfrac{685}{9}\approx76.111111"

"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum_iX_i)^2"

"=20930-\dfrac{408^2}{9}=2434"

"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum_iY_i)^2"

"=54191-\dfrac{685^2}{9}=2054.888889"

"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum_iX_i)(\sum_iY_i)"

"=32954-\dfrac{408(685)}{9}=1900.666667"

"slope=m=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{1900.666667}{2434}"

"=0.7809"

"n=\bar{Y}-m\bar{X}=76.111111-0.7809(45.333333)"

"=40.7111"

a.The regression equation is:


"Y=40.7111+0.7809X"

"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}"

"=\dfrac{1900.666667}{\sqrt{2434}\sqrt{2054.888889}}\approx0.849868"

Strong positive correlation.

The SY16C grades increases from "40.7111." The slope is "0.7809."


b.


"Y=40.7111+0.7809(60)"

"SY16C=88"